OpenStudy (anonymous):

How do you find dy/dx using implicit differentiation if cos(xy)=1+sin(y)

OpenStudy (anonymous):

\[-\sin(xy)\frac{ d }{ dx }(xy) = 0 + cosy \frac{ d }{ dx } (y)\]

OpenStudy (anonymous):

i used the chain rule above